OpenStudy (anonymous):

Help please! A tower in Centerville sends radio signal a certain distance (in miles) according to the equation: x^2 + y^2 = 3,600. The Martinez family lives 50 miles north and 70 miles east of Centerville. Can the Martinez family receive the radio signal from Centerville? In complete sentences, explain your reasoning.

OpenStudy (mathmate):

Calculate the square of distance from the tower, which is (50^2+70^2)=7400 Since 7400 > 3600, unfortunately the Martinez will not be able to make use of the tower in centerville, which requires the square of distance x^2+y^2 <=3600.

OpenStudy (anonymous):

Are you sure that's how you solve it? why would you square the 50 and 70?

OpenStudy (mathmate):

Usually we square the x- and y-coordinates, sum them, and take the square-root to find the distance. The given condition of x^2+y^2=3600 is equivalent to saying reception is possible within a radius of sqrt(3600)=60, which at first sight, the Martinez family will not get the reception because 70>60. To make the calculations easier to understand, I made the squares of the distances and compare the squares instead, which is verbatim with the requirements. To make things more complicated, the given condition says x^2+y^2=3600 (instead of <3600), which is not a reasonable requirement. Signal reception is usually expressed in an area within a given radius. I have assumed the more general interpretation. Hope my explanations are understandable.